Diffusion Morning Report 03/28/11

Ted Berg tberg at stanford.edu
Mon Mar 28 06:45:15 PDT 2011


Good Morning All,
     Looking good this morning. Tylan 1 controller is back and seems to 
be working well.TEOS pump should be back today. Tylan LTO  failed pump 
program. Thermco LTO needs a test after repairs. Thermco Poly 1 no dep. 
Amat Epi was back together on Friday not sure of status. Thermco 2 no 
TLC flow. No other major issues at this time.

					
	Date: 	3/28/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*yellow* 	*needs test was run over weekend* 	** 	** 	*3*
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*pump bad-out for repair* 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*20.67/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*9.6/30* 	** 	**
Tylan 11-LTO 	*yellow* 	*pump program failed* 	*3.28/30* 	** 	*1*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*down* 	*NO DEP* 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*yellow* 	*needs test after repair* 	** 	** 	**
AMAT Epi 	*yellow* 	*conditioning* 	** 	** 	*2*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 600 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110328/b1f34438/attachment.html>


More information about the diffusion mailing list